OpenStudy (anonymous):

John and Alan have a collection of x baseball cards. John has X/4 cards. What fraction of the cards does Alan have?

OpenStudy (anonymous):

So there are \(x\) total cards?

OpenStudy (anonymous):

And John has \((1/4)x\) of those cards...

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

So what happens when you subtract \(x - (1/4)x\)?

OpenStudy (anonymous):

\[ x-(1/4)x = [1-(1/4)]x \]

OpenStudy (anonymous):

Do you know how to subtract a fraction?

OpenStudy (anonymous):

sure dont

OpenStudy (anonymous):

First you need the denominators to be the same: \[ 1= \frac{4}{4} \]

OpenStudy (anonymous):

Now denominators are the same... \[ \frac44 - \frac14 \]The result has the same denominator (bottom)... the numerator (top) is the difference between numerators \[ \frac44 - \frac14 = \frac{4-1}4 = \frac34 \]

OpenStudy (anonymous):

So Alan has \((3/4)\) of the total cards... meaning \((3/4)x\) cards.

OpenStudy (anonymous):

i already picked that answer, and I got it wrong. so the choices are between A. x/3 B. 3x/4 and D. 3x

OpenStudy (anonymous):

I'm thinking it could be 3x/4 though

OpenStudy (anonymous):

yes you should know by now it is that one.